Evaluate the following line integrals along with the curve C.

(a)

C (x2 − 2y2 ) ds; C is the line segment parametrized by r(t) = < t √ 2 , t √ 2> for 0 ≤ t ≤ 4.


1
Expert's answer
2021-12-29T17:12:20-0500

"\begin{aligned}
& \int_{c}\left(x^{2}-2 y^{2}\right) d s, r(t)=\left<\frac{1}{\sqrt{2}}, \frac{t}{\sqrt{2}}\right> \text { for } 0 \leq t \leq 4\\
&F(x, y)=x^{2}-2 y^{2}\\
&F(x+t))=\left(\frac{t}{\sqrt{2}}\right)^{2}-2\left(\frac{t}{\sqrt{2}}\right)^{2}\\
&=\frac{t^{2}}{2}-t^{2}=-\frac{t^{2}}{2}\\
&r^{\prime}(t)=\left(\frac{1}{\sqrt{2}}, \frac{1}{\sqrt{2}}\right)\\
&\int_{0}^{4}\left(\frac{-t^{2}}{2}\right) \cdot\left(\frac{1}{\sqrt{2}}, \frac{1}{\sqrt{2}}\right) \cdot d t\\
&\int_{0}^{4} 2\times \frac{1}{\sqrt{2}}\left(-\frac{t^{2}}{2}\right) d t\\
&\left.\frac{1}{\sqrt{2}} \int \frac{-t^{3}}{3}\right|_{0} ^{4}=\frac{1}{\sqrt{2}} \frac{(-4)^{3}}{3}=-\frac{64}{\sqrt{2} \times 3}\\
&=\frac{-64 \sqrt{2}}{6}=\frac{-32}{3} \sqrt{2}
\end{aligned}"